# The Machine Learning Control Method for Counterfactual $*{}^{*}$

### v1

@@ 3.1 Departures from the standard machine learning approach
The standard ML approach is to randomly split the sample into two sets, containing, for instance, 2/3 and 1/3 of observations. One then uses the first set to train ML algorithms (training set) and the second to test them (testing set). This introduces a “firewall” principle: none of the data involved in generating the prediction function is used to evaluate it (Mullainathan and Spiess,, 2017). The out-of-sample performance of the model on the unseen (held-out) data of the testing set can be considered a reliable measure of the “true” performance on future data. In order to solve the bias-variance trade-off and prevent overfitting, one can rely on automatic tuning using tools such as random k-fold CV on the training sample to select the best-performing values of the tuning parameters in terms of an a priori defined metric, such as the Mean Squared Error (MSE).
@@
Second, and related, we do not carry out hyperparameter tuning and model selection with random k-fold CV. The panel dimension, in fact, creates an additional challenge regarding how to exactly implement CV, because standard CV does not account for the temporal structure of the data (Arkhangelsky and Imbens,, 2023). To address this challenge, we introduce a resampling technique suited for forecasting tasks on panel data—panel CV—which is fully described below.
Finally, a key concern in ML regards the trade-off between accuracy and interpretability. Such a trade-off is relevant when ML is used for tasks that take into consideration transparency and interpretability aspects. In the case of the MLCM, we argue that preserving transparency is important because higher interpretability of the estimated counterfactuals bolsters the credibility of the proposed approach (Abadie,, 2021). Since our method can be used with any supervised ML routine, users should decide on the basis of a comparative performance assessment across a mix of models characterized by different layers of complexity by carefully balancing any improvements in performance from complex models against the loss of interpretability that comes with their use.
@@ 3.2 Implementation
The implementation of the MLCM requires ten empirical steps, divided into the Design Stage and the Analysis Stage. The full process is summarized in Box 3.2 and described in detail below.
Preliminary   0. Data splitting. Split the full sample based on the treatment date: employ only pre-treatment data throughout the Design Stage; use post-treatment data only for estimating causal effects in the Analysis Stage.   Design Stage   1. Algorithm selection. Select one or more supervised ML algorithms. 2. Principled input selection. Build a large initial dataset on the basis of domain knowledge. To maximize forecasting performances, it is possible to use feature engineering, feature selection, and heuristic rules to pre-process the data and select a subsample of the most relevant predictors. 3. Panel cross-validation. For each selected algorithm, tune hyperparameters via panel CV (see Figure 1 for the case of one-step ahead forecasts). 4. Performance assessment. Assess average performance metrics (e.g., MSE) for all the selected algorithms and check what is the best-performing version of the MLCM. 5. Diagnostic and placebo tests. Implement a battery of diagnostic and placebo tests to bolster the credibility of the research design.   Analysis Stage             6. Final model selection. On the basis of the comparative performance assessment in the Design Stage, pick the best-performing model and use that for the Analysis Stage. Start by re-training the model on the full pre-treatment sample using the hyperparameter(s) selected in the Design Stage. 7. Counterfactual forecasting. For each unit $i$, forecast the post-treatment counterfactual outcome $\widehat{\operatorname{Y}}_{i,t_{0}+k|t_{0}}$. In case of a large-scale shock affecting important covariates, either rely solely on their past lags or repeat steps 1–6 to forecast $\widehat{\mathbf{X}}_{i,t_{0}+k|t_{0}}(0)$ and use these values to improve the forecast of $\widehat{\operatorname{Y}}_{i,t_{0}+k|t_{0}}$. 8. Estimation of treatment effects. For each unit $i$, estimate the individual treatment effect in Equation (4) by taking the difference between the observed post-treatment outcome $\operatorname{Y}_{i,t_{0}+k}$ and the ML-generated potential outcome $\widehat{\operatorname{Y}}_{i,t_{0}+k|t_{0}}$. Estimate other causal estimands, such as the ATE from Equation (5) or the temporal ATE from Equation (7), by aggregating the individual estimates. 9. Treatment effect heterogeneity. To uncover heterogeneity, data-driven CATEs can be estimated as in Equation (6) via a regression tree analysis with the individual treatment effects as the outcome variable and a set of exogenous predictors potentially associated with treatment effect heterogeneity. 10. Inference. Compute standard errors for the ATE and CATEs via block-bootstrap.
@@
In the Design Stage, the first step involves the selection of supervised ML algorithms that will play the horse-race of performance testing on the pre-treatment data. [^17] It is possible to stack several different ML algorithms and form complex ensemble learners, but we refrain from their use because, as stated above, we deem it important to retain some degree of transparency in counterfactual building. In the second step, we recommend deploying some tweaks involving feature pre-selection and engineering drawing from consolidated practices in applied predictive modeling (Kuhn and Johnson,, 2013). A crucial aspect of this pre-processing is the strategy for selecting predictors. There are two main schools of thought: those who advocate for purely data-driven selection argue that one should build a dataset as large as possible, and then let the algorithm autonomously decide which variables matter for the forecasting task. Others stress the importance of subject matter knowledge: the researcher should select ex ante the relevant predictors, and then feed only those to the algorithm. The underlying rationale is that subject matter knowledge can separate meaningful from irrelevant information, eliminating detrimental noise and enhancing the underlying signal (Kuhn and Johnson,, 2013). We propose a hybrid approach: build a large initial dataset on the basis of domain knowledge, then adopt preliminary and data-driven variable selection criteria to drop non-informative predictors.
@@
As outlined in the causal framework, counterfactual forecasting is carried out by using an information set mainly comprising lagged values of outcomes and covariates. However, determining the optimal number of lags to include is an empirical question. We recommend including at least two lagged values of both outcomes and covariates, and then using a data-driven approach to select a subsample of the most relevant features. The latter step allows for a reduction of the risk of overfitting and degradation of forecasting performances as well as facilitating interpretability. Finally, remember that feature engineering and data pre-processing matter too because how the predictors enter into the model is also important (Kuhn and Johnson,, 2013).
To carry out model selection and validation on the pre-treatment sample, we propose a panel CV approach. In our setting, using an alternative CV procedure is necessary since ML methods do not natively handle longitudinal data and are designed for predicting rather than forecasting. Our panel CV approach adapts time series CV based on expanding training windows to a panel setting. [^18] Here we use an expanding window (as we are in a short-panel setting), but the procedure can also be implemented using a rolling window approach, which might be more appropriate in specific cases (e.g., with long panels or non-stationarity of the outcome). The intuition—for the case of one-step ahead forecasts, but the routine is easily adapted to multi-step ahead forecasts—is provided in Figure 1 and constitutes an adaptation from Hyndman and Athanasopoulos, (2021).
@@
Figure 1: Panel cross-validation (one-step ahead forecasting). Note: this procedure is carried out using exclusively pre-treatment data. Light gray observations form the training sets; dark gray ones constitute the test sets. All the training windows also include past outcomes in the set of predictors.
In short, we establish cutoff points in the temporal dimension as $\{t_{0}-s,t_{0}-s+1,\dots,t_{0}-1\}$, with $s$ denoting a strictly positive integer such that $t_{0}-s\geq 1$. At the first CV step, these cutoffs delineate a training set $\mathcal{T}_{1}=\{\operatorname{Y}_{i,t}:t=1,\dots,t_{0}-s\}$ and a validation set $\mathcal{V}_{1}=\{\operatorname{Y}_{i,t}:t=t_{0}-s+1\}$. At the second CV step, the training set becomes $\mathcal{T}_{2}=\{\operatorname{Y}_{i,t}:t=1,\dots,t_{0}-s+1\}$ and the validation set $\mathcal{V}_{2}=\{\operatorname{Y}_{i,t}:t=t_{0}-s+2\}$ and so on, until the final CV step, where the validation set includes the observations in the last time period before the intervention. This sequential procedure ensures that, for each unit, there are no “future” observations in the training set and no “past” observations in the validation set. At each step, the parameters’ values yielding the best predictive performance in the validation step (e.g., those minimizing the MSE) are stored and then averaged across all CV steps. This also provides summary measures of average model performance on the pre-treatment data, which can be screened to select the winner of the horse-race.
@@
To provide evidence about internal validity, we suggest running diagnostic checks (e.g., showing that the distribution of the pre-treatment forecasting errors is approximately Gaussian and centered around zero) as well as placebo tests, which have become a key device for assessing the credibility of research designs in observational settings (Eggers et al.,, 2021). Following Liu et al., (2022), panel placebo tests can be implemented by hiding one or more periods of observations right before the onset of the treatment and using a model trained on the rest of the pre-treatment periods to predict the untreated outcomes of the held-out period(s). If the identifying assumptions are valid, the differences between the observed and forecast outcomes in those periods should be close to zero. Importantly, given that we estimate unit-level treatment effects, we are able to do more than just show that the average differences are close to zero (the standard practice in most event-study designs) and test whether most unit-level placebo differences are close to zero. The Design Stage thus ends with a battery of performance, diagnostic, and placebo tests.
@@
When dealing with a single post-intervention period, step 7 of the Analysis Stage is straightforward. However, when performing a multi-step ahead forecast using a data-driven ML routine, one must take special care to avoid including post-treatment outcomes in the prediction. To achieve this, we adopt the following procedure: we first forecast the counterfactual outcome absent the policy at time $t_{0}+1$ (by relying only on pre-treatment information) and obtain an estimate of $\widehat{\operatorname{Y}}_{i,t_{0}+1|t_{0}}$. Then we proceed one step ahead, by including $\widehat{\operatorname{Y}}_{i,t_{0}+1|t_{0}}$ in the past information set and using it to forecast $\widehat{\operatorname{Y}}_{i,t_{0}+2|t_{0}}$. We recursively substitute the forecasted value until we reach the last time point, $\widehat{\operatorname{Y}}_{i,t_{0}+T|t_{0}}$. [^19] This is analogous to what is typically done in the usual time series setting. Consider for example an AR(1) process: $y_{t}=\phi y_{t-1}+\epsilon_{t}$, with $||\phi||<1$. The 1-step ahead forecast is given by $y_{t+1|t}=\operatorname{E}[y_{t+1}|\mathcal{I}_{t}]=\phi\operatorname{E}[y_{t}|\mathcal{I}_{t}]=\phi y_{t}$; the 2-step ahead forecast is readily computed from the forecasted value $y_{t+1|t}$, i.e., $y_{t+2|t}=\operatorname{E}[y_{t+2}|\mathcal{I}_{t}]=\phi\operatorname{E}[y_{t+1}|\mathcal{I}_{t}]=\phi^{2}y_{t}$. Furthermore, note that the same procedure can be applied to forecast post-treatment values of important covariates affected by the intervention. Like in the case of past outcomes (see Section 2.4), there are two options for handling covariates that could be impacted by the intervention: either we impose that $f(\cdot)$ contains the past $q$ lags of $\mathbf{X}_{i,t_{0}+k}$ up to $t_{0}$, or we assume that such covariates follow Equation (2) and use the ML algorithms and the panel CV routine to forecast counterfactual values of $\widehat{\mathbf{X}}_{i,t_{0}+k|t_{0}}(0)$, which we can then incorporate into the prediction of $\widehat{\operatorname{Y}}_{i,t_{0}+k|t_{0}}$.
@@
Next, data-driven CATEs are computed via a regression tree analysis on the full sample of estimated treatment effects. Specifically, this approach uses the estimated treatment effects as the outcome variable, regresses them on many potentially associated variables, and lets the algorithm pick the main predictors and their critical thresholds. The resulting tree reports the group-average treatment effects for all units in each terminal node. As for causal trees and forests (Athey and Imbens,, 2016; Wager and Athey,, 2018), this data-driven search for heterogeneity of causal effects removes a major degree of discretion because the researcher can only select the set of covariates that can be used by the tree to build the subgroups. However, the two approaches differ regarding both purpose and implementation: our data-driven technique is a post-estimation approach aimed at automatically recovering and visualizing the relevant heterogeneity dimensions, while causal trees and forests are counterfactual methods for the direct estimation of heterogeneous treatment effects, which they achieve by leveraging control units. Moreover, we are not interested in the out-of-sample performance of the regression tree, but in retrieving CATEs for the entire population of interest. To this end, there is no need either to split the sample into training and testing sets or to prune the tree by adjusting the complexity parameter.
@@
Finally, standard errors and confidence intervals for the ATE and CATEs are estimated through the block-bootstrap approach described in Appendix A.

### v2

@@ 3.1 Departures from the standard machine learning approach
The standard ML approach is to randomly split the sample into two sets, containing, for instance, 2/3 and 1/3 of observations. One then uses the first set to train ML algorithms (training set) and the second to test them (testing set). This introduces a “firewall” principle: none of the data involved in generating the prediction function is used to evaluate it (Mullainathan and Spiess, 2017). The out-of-sample performance of the model on the unseen (held-out) data of the testing set can be considered a reliable measure of the “true” performance on future data. In order to solve the bias-variance trade-off and prevent overfitting, one can rely on automatic tuning using tools such as random k-fold CV on the training sample to select the best-performing values of the tuning parameters in terms of an a priori defined metric, such as the MSE.
@@
Second, and related, we do not carry out hyperparameter tuning and model selection with random k-fold CV. The panel dimension creates an additional challenge regarding how to implement CV, because standard CV does not account for the temporal structure of the data (Arkhangelsky and Imbens, 2024). To address this challenge, we introduce a resampling technique suited for forecasting tasks on panel data—panel CV—which is described below.
Finally, a key concern in ML regards the trade-off between accuracy and interpretability. Such a trade-off is relevant when ML is used for tasks that take into consideration transparency aspects. In principle, our method can be used with any supervised ML routine, including black-box techniques like deep neural networks. However, maintaining some degree of transparency can be important, because higher interpretability of the estimated counterfactuals bolsters the credibility of the proposed approach (Abadie, 2021). In practice, users should decide on the basis of a comparative assessment across a mix of models characterized by different layers of complexity by balancing any improvements in performance from complex models against the loss of interpretability associated with their use.
@@ 3.2 Implementation
The MLCM is rooted in the idea of developing a ML forecasting model that can closely reproduce the outcome trajectories of treated units in the pre-intervention window, so that any post-intervention divergence between the observed and forecasted outcomes can be attributed to the treatment under the identification assumptions. To achieve this, the implementation of the MLCM requires ten empirical steps, which are divided into the Design Stage and the Analysis Stage. The full process is summarized in Box 3.2 and described in detail below.
In the Design Stage, the first step involves the selection of supervised ML algorithms that will play the horse-race of performance testing on the pre-treatment data. This competition among multiple methods is conceptually analogous to the way the best ML learner is selected in the double/debiased ML method (Chernozhukov et al., 2018). [^16] The MLCM can leverage any supervised ML algorithm, including deep learning techniques, some of which, such as RNNs and transformers, are explicitly designed to learn temporal dynamics; however, they cannot be employed in the panel settings we focus on (small T, large N), as they require a large number of time periods to be applicable. It is also possible to stack several different ML algorithms and form complex ensemble learners, but this would come at the cost of a substantial loss in transparency.
@@
In the second step, we recommend deploying some tweaks involving feature pre-selection and engineering drawing from consolidated practices in applied predictive modeling (Kuhn and Johnson, 2013). A crucial aspect of this pre-processing is the strategy for selecting predictors. There are two main schools of thought: those who advocate for purely data-driven selection argue that one should build a dataset as large as possible, and then let the algorithm autonomously decide which variables matter for the prediction task. Others stress the importance of subject matter knowledge: the researcher should select ex ante the relevant predictors, and then feed only those to the algorithm. The rationale is that subject matter knowledge can separate meaningful from irrelevant information, eliminating detrimental noise and enhancing the underlying signal (Kuhn and Johnson, 2013). We propose a hybrid approach: build a large initial dataset on the basis of domain knowledge, then adopt preliminary and data-driven variable selection criteria to drop non-informative predictors.
@@
As outlined in the causal framework, counterfactual forecasting is carried out by using an information set mainly comprising lagged values of outcomes and covariates. However, determining the optimal number of lags to include is an empirical question. We recommend including at least two lagged values of both outcomes and covariates, and then using a data-driven approach to select a subsample of the most relevant features. The latter step allows for a reduction of the risk of overfitting and degradation of performances as well as facilitating interpretability. Finally, feature engineering and data pre-processing matter too because how the predictors enter into the model is also important (Kuhn and Johnson, 2013).
To carry out model selection and validation on the pre-treatment sample, we propose a panel CV approach. Using an alternative CV procedure is necessary since ML methods do not natively handle longitudinal data and are designed for predicting rather than forecasting.
{mdframed}
[ innerlinewidth=0.5pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin = 10pt, innerbottommargin=10pt, frametitle=                                                      Box 3.2: MLCM implementation, frametitlerule=true, frametitlerulewidth=1pt]
Preliminary
  1. 0.
    Data splitting. Split the full sample based on the treatment date: employ only pre-treatment data throughout the Design Stage; use post-treatment data only for estimating causal effects in the Analysis Stage.
Design Stage
  1. 1.
    Algorithm selection. Select one or more supervised ML algorithms.
  2. 2.
    Principled input selection. Build a large initial dataset on the basis of domain knowledge. To maximize forecasting performances, it is possible to use feature engineering, feature selection, and heuristic rules to pre-process the data and select a subsample of the most relevant predictors.
  3. 3.
    Panel cross-validation. For each selected algorithm, tune hyperparameters via panel CV (see Figure 1 and Algorithms 1 and 2).
  4. 4.
    Performance assessment. Assess average performance metrics (e.g., MSE) for all the selected algorithms and check what is the best-performing version of the MLCM.
  5. 5.
    Diagnostic and placebo tests. Implement a battery of diagnostic and placebo tests to bolster the credibility of the research design.
    Analysis Stage
  6. 6.
    Final model selection. On the basis of the comparative performance assessment in the Design Stage, pick the best-performing model and use that for the Analysis Stage. Start by re-training the model on the full pre-treatment sample using the hyperparameter(s) selected in the Design Stage.
  7. 7.
    Counterfactual forecasting. For each unit $i$, forecast the post-treatment counterfactual outcome $\widehat{\operatorname{Y}}_{i,t_{0}+k|t_{0}}(0)$. In case of a large-scale shock affecting important covariates, either rely solely on their past lags or repeat steps 1–6 to forecast $\widehat{\mathbf{X}}_{i,t_{0}+k|t_{0}}(0)$ and use these values to improve the forecast of $\widehat{\operatorname{Y}}_{i,t_{0}+k|t_{0}}(0)$. Leverage Explainable Artificial Intelligence tools to enhance the model’s explainability and transparency of the estimated counterfactual.
  8. 8.
    Estimation of treatment effects. For each unit $i$, estimate the individual treatment effect as in Equation (6) by taking the difference between the observed post-treatment outcome $\operatorname{Y}_{i,t_{0}+1}$ and the ML-generated potential outcome $\widehat{\operatorname{Y}}_{i,t_{0}+1|t_{0}}(0)$. For $k\geq 2$, if the selected ML algorithm in Step 4 of the Design Stage is linear, estimate $\hat{\tau}_{t_{0}+k}$ as in Equation (10), otherwise use Equation (12). Estimate other causal estimands, such as the ATE from Equation (13) or the temporal ATE from Equation (14), by aggregating the individual estimates.
  9. 9.
    Treatment effect heterogeneity. To uncover heterogeneity, data-driven CATEs can be estimated as in Equation (13) via a regression tree analysis with the individual treatment effects as the outcome variable and a set of variables potentially associated with treatment effect heterogeneity.
  10. 10.
    Inference. Compute standard errors for the ATE and CATEs via block-bootstrap.
@@
Figure 1: Panel cross-validation (one-step-ahead forecasting). Notes: this procedure is carried out using exclusively pre-treatment data. Light gray observations form the training sets; dark gray ones constitute the test sets. All the training windows also include past outcomes in the set of predictors.
Our panel CV approach adapts time series CV based on expanding training windows to a panel setting. [^17] Here we use an expanding window (as we are in a short-panel setting), but the procedure can also be implemented using a rolling window approach, which might be more appropriate in specific cases (e.g., with long panels or non-stationarity of the outcome). The intuition is provided in Figure 1 and constitutes an adaptation from Hyndman and Athanasopoulos (2021).
We start from a set of candidate algorithms $\mathcal{H}=(h_{1},\dots,h_{J})$, each having its own set of parameters $\theta^{(j)}$ and hyperparameters $\gamma^{(j)}$. For example, in the empirical application, we use four learners: LASSO, PLS, random forest and stochastic gradient boosting. For each learner, we carry out panel CV as detailed in Algorithm 1. In short, the ML algorithms are trained repeatedly using an expanding window approach, with hyperparameters tuned at each step to minimize forecast errors. This sequential procedure ensures that, for each unit, there are no “future” observations in the training set and no “past” observations in the validation set. Note that the same procedure can be applied to forecast post-treatment values of important covariates affected by the intervention. We can assume that such covariates follow Equation (2) and use the ML algorithms and the panel CV routine to forecast counterfactual values of $\widehat{\mathbf{X}}_{i,t_{0}+k|t_{0}}(0)$, which we can then incorporate into the prediction of $\widehat{\operatorname{Y}}_{i,t_{0}+k|t_{0}}$ as in Liu et al. (2020).
0:  $\mathcal{H}=(h_{1},\dots,h_{J})$, $t_{0}$
1:  for $j$ in $1:J$, where $J$ is the total number of candidate algorithms do
2:     for $s$ in $11:t_{0}-$ 1 do
3:        Split the data into a training set $\mathcal{T}=\{\operatorname{Y}_{i,t},Z_{i,t}\}_{t=1}^{s}$ and a validation set $\mathcal{V}=\{\operatorname{Y}_{i,t},Z_{i,t}\}_{t=s+1}$, where $Z_{i,t}=\{\operatorname{Y}_{i,t-1}^{(p)},\mathbf{X}^{(q)}_{i,t}\}$ is a set of features including past lags of the outcome and covariates that should be related to $\operatorname{Y}_{i,t}$.
4:        Fit the model in the training set by optimizing a performance measure, e.g., minimizing MSE,
$\hat{\theta}^{(j)}=\operatorname*{argmin}_{\theta}{\sum_{i,t\in\mathcal{T}}(\operatorname{Y}_{i,t}-h_{j}(Z_{i,t};\gamma^{(j)},\theta^{(j)}))^{2}}$
and compute the $1$-step-ahead forecast, i.e., $\operatorname{E}[\operatorname{Y}_{i,s+1}|Z_{i,s+1}]=h_{j}(Z_{i,s+1};\gamma^{(j)},\hat{\theta}^{(j)})$.
5:        Tune hyperparameters in the validation set by optimizing the one-step-ahead forecast error,
$\hat{\gamma}^{(j)}=\operatorname*{argmin}_{\gamma}\sum_{i}(\operatorname{Y}_{i,s+1}-h_{j}(Z_{i,s+1};\gamma^{(j)},\hat{\theta}^{(j)}))^{2}$
6:        Repeat steps 3-5 for all $s$ up to $t_{0}-1$, the optimal values for the hyperparameters will be those that minimize the forecast errors across all horizons.
7:     end for
8:     Repeat steps 1-6 for all $j$ and choose the algorithm in $\mathcal{H}$ which minimizes the one-step-ahead forecast error.
9:  end for
Algorithm 1 Panel CV for one-step-ahead forecasting
@@
To provide evidence about internal validity, we suggest running diagnostic checks (e.g., showing that the distribution of the pre-treatment forecasting errors is approximately Gaussian and centered around zero) as well as placebo tests, which have become a key device for assessing the credibility of research designs in observational settings (Eggers et al., 2024). Following Liu et al. (2022), panel placebo tests can be implemented by hiding one or more periods of observations right before the onset of the treatment and using a model trained on the rest of the pre-treatment periods to predict the untreated outcomes of the held-out period(s). If the identifying assumptions are valid, the differences between the observed and forecast outcomes in those periods should be close to zero. [^18] If large discrepancies, i.e., forecasting errors, are found at this stage, this would indicate pre-intervention shocks or policy changes occurring between pre-treatment periods. These factors should be accounted for in the model to bridge the gap between observed and forecasted pre-treatment outcomes. Given that we estimate unit-level treatment effects, we are also able to test whether most unit-level placebo differences are close to zero. The Design Stage thus ends with a battery of performance, diagnostic, and placebo tests.
@@
When dealing with a single post-intervention period, step 7 of the Analysis Stage is straightforward. However, when performing a multi-step-ahead forecast using a data-driven ML routine, one must take special care to avoid including post-treatment outcomes in the prediction. Multi-step-ahead forecast becomes more challenging when the selected ML algorithm is non-linear, as we cannot simply plug-in the forecasted value $\widehat{\operatorname{Y}}_{t_{0}+1|t_{0}}(0)$ into the already estimated model. As explained in Section 2.3, one solution is to re-estimate the potential outcome model for values of $k\geq 2$. Importantly, at the end of the panel CV procedure, we know the selected ML algorithm and can adjust our forecasting strategy accordingly. Specifically, we propose using Equation (3) to model post-intervention potential outcomes, as this approach has proven effective in identifying non-linear, multi-step-ahead causal effects, and estimate them with Equation (12). From a practical standpoint, this requires repeating the panel CV for each step of the forecasting horizon. The detailed procedure is described in Algorithm 2.
0:  $h^{win}$ (selected model after Algorithm 1).
1:  Compute $\widehat{\operatorname{Y}}_{i,t_{0}+1|t_{0}}=\widehat{h}^{win}(y_{i,t_{0}}^{(p)},\mathbf{x}_{i,t_{0}+1}^{(q)})$
2:  if $h^{win}$ is a non-linear model and the forecasting horizon $k\geq 2$ then
3:     for $k=2$ do
4:        Start from a set of candidate algorithms $\mathcal{G}=(g_{2}^{(1)},\dots,g_{2}^{(j)})$
5:        Split the data into a training set $\mathcal{T}=\{\operatorname{Y}_{i,t_{0}}^{(p)},\mathbf{X}_{i,t_{0}}^{(q)}\}$ and a validation set $\mathcal{V}=\{\widehat{\operatorname{Y}}_{i,t_{0}+1|t_{0}},\mathbf{X}_{i,t_{0}+1}^{(q)}\}$.
6:        Fit the model $g_{2}^{(j)}$ in the training set by optimizing a performance measure, e.g., minimizing MSE,
$\hat{\theta}^{(j)}=\operatorname*{argmin}_{\theta}{\sum_{i,t_{0}\in\mathcal{T}}(\operatorname{Y}_{i,t_{0}}-g_{2}^{(j)}(\operatorname{Y}_{t_{0}-1}^{(p)},\mathbf{X}_{i,t_{0}}^{(q)};\gamma^{(j)},\theta^{(j)}))^{2}}$
and compute the $1$-step-ahead forecast based on $g_{2}$, i.e., $g_{2}^{(j)}(\operatorname{Y}_{t_{0}}^{(p)},\mathbf{X}_{i,t_{0}+1}^{(q)};\gamma^{(j)},\hat{\theta}^{(j)})$.
7:        Tune hyperparameters in the validation set by optimizing the one-step-ahead forecasting error,
$\hat{\gamma}^{(j)}=\operatorname*{argmin}_{\gamma}\sum_{i}(\widehat{\operatorname{Y}}_{i,t_{0}+1|t_{0}}-g_{2}^{(j)}(\operatorname{Y}_{t_{0}}^{(p)},\mathbf{X}_{i,t_{0}+1}^{(q)};\gamma^{(j)},\hat{\theta}^{(j)}))^{2}.$
8:        Choose the candidate algorithm in $\mathcal{G}$ which minimizes the one-step-ahead forecast error, retrain the model including the validation set and compute the one-step-ahead forecast $\widehat{\operatorname{Y}}_{t_{0}+2|t_{0}}=\widehat{g}_{2}(\widehat{\operatorname{Y}}_{t_{0}+1|t_{0}},\operatorname{Y}_{t_{0}}^{(p)},\mathbf{X}_{t_{0}+2}^{(q)})$.
9:        Repeat steps 4-8 for all $k>2$
10:     end for
11:  end if
Algorithm 2 Panel CV for non-linear multi-step-ahead forecasting
During the counterfactual forecasting step, especially when opting for more complex and less transparent techniques, users should consider leveraging innovations in Interpretable Machine Learning and Explainable Artificial Intelligence (Molnar, 2020), such as model-agnostic Shapley Additive Explanations (SHAP) (Lundberg and Lee, 2017), to enhance explainability and transparency of the counterfactual building process (Abadie, 2021).
@@
After estimating individual effects, data-driven CATEs can be computed via a regression tree analysis. Specifically, this approach uses the estimated treatment effects as the outcome variable, regresses them on many potentially associated variables, and lets the algorithm pick the main predictors and their critical thresholds. The resulting tree reports the group-average treatment effects for all units in each terminal node. As for causal trees and forests (Athey and Imbens, 2016; Wager and Athey, 2018), this data-driven search for heterogeneity of causal effects removes a major degree of discretion because the researcher can only select the set of covariates that can be used by the tree to build the subgroups. However, the two approaches differ regarding both purpose and implementation: our data-driven technique is a post-estimation approach aimed at automatically recovering and visualizing the relevant heterogeneity dimensions, while causal trees and forests are counterfactual methods for the direct estimation of heterogeneous treatment effects, which they achieve by leveraging control units. Moreover, we are not interested in the out-of-sample performance of the regression tree, but in retrieving CATEs for the entire population of interest. To this end, there is no need to split the sample into training and testing sets or to prune the tree by adjusting the complexity parameter. However, it may be necessary to set a pre-specified minimum node size to preserve the interpretability of the resulting tree.
@@
Finally, standard errors and confidence intervals for the ATE and CATEs are estimated through the block-bootstrap approach described in Supplemental Appendix C. [^19] In staggered adoption settings, uncertainty quantification with block-bootstrapping can be extended to other estimands, such as group-time average treatment effects. See the replication exercise in Supplemental Appendix F.